OpenStudy (anonymous):

The steps below show the incomplete solution to find the value of x for the equation 5x –2x –3 = –2 + 15: Step 1: 5x –2x –3 = –2 + 15 Step 2: 5x –2x –3 = 13 Step 3: 3x –3 = 13 Which of these is most likely the next step? 3x = 16 3x = 10 3x = 39 3x = 3

OpenStudy (anonymous):

after step number three you want to have 3x by itself so you need to move the -3.

OpenStudy (anonymous):

please don't give direct answers @Hannah_Waller

OpenStudy (anonymous):

okay so anyways you need to move the -3. to do that you add 3 to both sides. 13+3 is ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

yep so now it looks like this: 3x=16

OpenStudy (anonymous):

The steps below show the incomplete solution to find the value of x for the equation 5x –2x –3 = –2 + 15: Step 1: 5x –2x –3 = –2 + 15 Step 2: 5x –2x –3 = 13 Step 3: 3x –3 = 13 Continue..... Step 4: 3x=13+3 Step 5: 3x=16 Step 6: x=16/3

OpenStudy (anonymous):

wow good job @njtwins !!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\(\huge\cal\color{red}{Welcome~}\color{blue}{to~}\color{green}{Open}\color{orange}{Study!}\)

OpenStudy (anonymous):

thanks for the help guys! i can only medal the best helper!

OpenStudy (anonymous):

@njtwins helped the most so medal him!

OpenStudy (anonymous):

who?